### v1

@@ 1 Introduction
Motivated by this observation, we propose LayoutFlow, a generative framework based on Flow Matching [32, 33, 1, 44] that can produce high-quality layouts while requiring less time than previous methods. Flow Matching has recently been introduced as a powerful generative framework and demonstrated strong performance on various tasks including image generation [34]. Intuitively, flow-based models try to learn a flow that moves samples from a base distribution to a target distribution usually defined by the training data. In this work, we investigate how to apply Flow Matching for layout generation and show that flows offer a more intuitive, geometrical interpretation of the generation process compared to previous models.
@@
In the context of layout generation, the differences between diffusion-based and flow-based models become apparent when visualizing the trajectories created by the generation process, as illustrated in Fig. 1. The trajectories show how the sample from the base distribution is moved by the flow or backward diffusion process until it reaches the final layout prediction. While our LayoutFlow model produces smooth trajectories with short path lengths, diffusion-based models follow a noisy trajectory that constantly changes directions due to the added noise. As a result, diffusion models tend to require more sampling steps, i.e. longer inference time, compared to flow-based models. Altogether, we argue that learning how to move the initial sample straight toward the final prediction is much more natural than trying to do so under additional noise. Overall, our contributions can be summarized as follows:
@@
    On top of generating simpler trajectories, flow-based models generally also offer more choices than diffusion models, such as different prior distributions or training trajectories. We therefore extensively explore this additional flexibility offered by Flow Matching in the context of layout generation.
@@ 2.1 Layout Generation
Over the years, layout generation has been investigated using various methods. Early works started exploring how to generate layouts by minimizing an energy function based on pre-defined constraints [37, 36]. With the advent of generative machine learning techniques, however, research shifted to a more data-centric learning approach and started to use Variational Autoencoders (VAE) [47, 13, 2, 27], and Generative Adversarial Networks (GAN) [29, 30, 25]. While earlier works explored a variety of different architectures, such as Graph Convolutional Networks [27], Long-Short Term Memories (LSTM) [13, 24] or Transformers [47, 45, 14], most recent research has converged to using a Transformer-based architecture due to its flexibility and strong performance.
Diffusion-based Models. As diffusion models have proven to generate more diverse data and are generally easy to train, especially compared to GANs, most recent research has shifted towards exploring the diffusion process for layout generation tasks. Since layouts are represented by categorical data defining types of elements and continuous numerical values describing the element position and size, applying classical diffusion models is not straightforward. As a result, different approaches have incorporated a diffusion loss into their training process, which can be divided into continuous [18] and discrete diffusion models [3]. For the discrete case, the continuous coordinates are quantized and interpreted as different states along with the categorical element type [22, 20, 48, 15]. In this discrete scenario, the forward diffusion process is modeled as a random walk between discrete states that in addition to the quantized state also include a mask state, which removes elements from the canvas. While LayoutDM [22] and LDGM [20] rely on the mask-and-replace strategy proposed in [12], LayoutDiffusion [48] introduces a new mild forward process that is closer to the continuous process while still increasingly masking elements over time.
@@ 2.2 Flow Matching and Diffusion Models
Diffusion models, which were first proposed by Sohl-Dickenstein et al. [40] and later popularized by Ho et al. [18], can be formulated similarly to flow-based models. In fact, the term Flow Matching was inspired by the similar Score Matching loss [21] used by diffusion models. However, flows are characterized by an ODE, whereas the diffusion process can be formulated as an SDE with an additional stochastic component described by Brownian Motion [42]. As a result, diffusion models are restricted to a Gaussian as a base distribution, while flow-based models can freely choose the base distribution, which we also explore in our experiments. Notably, Song et al. [41] proposed DDIM, an alternative sampling method that makes it possible to sample using an ODE formulation on a model trained with an SDE forward process, effectively leading to smoother trajectories. In the context of layout generation, however, models have been using the SDE formulation following DDPM [18]. Our argument for introducing Flow Matching instead of just using DDIM for smoother trajectories lies in its increased flexibility and simplicity. Moreover, if the goal is to sample using an ODE, it seems more natural to also use an ODE during training, instead of the more complex SDE used in diffusion.
@@ 3 Preliminary: Flow Matching
In this section, we first present the mathematical definition of a flow and then introduce how to train a neural network to estimate a flow via Flow Matching. The term flow in Flow Matching refers to a mapping between samples of two distributions. In general, flow-based models aim to estimate the flow between a known source probability distribution $p_{0}(x)$, e.g. a Gaussian distribution, and the typically more complex data distribution $p_{1}(x)$. Given a data point $x\in\mathbb{R}^{d}$, the flow $\phi:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ can be defined by the following Ordinary Differential Equation (ODE)
@@
However, this loss alone does not allow us to train the model as we usually do not have access to $p_{t}$ nor $v_{t}$. As an extension, Lipman et al. [32] have proposed a Conditional Flow Matching objective, which they show to be equal up to a constant independent of $\theta$, and is given as
@@
with $v_{t}(x|z)$ denoting a conditional vector field generating the conditional probability path $p_{t}(x|z)$ and $q(z)$ being a distribution over a latent conditioning variable $z$. By introducing the conditional formulation of the Flow Matching objective, training a neural network becomes possible by choosing and training with a conditional vector field and associated probability path. For an overview of the various choices and more details on Conditional Flow Matching, we recommend [32, 44].
@@ 4.1 Data Representation
where we split each element into its geometrical information $\mathbf{g}^{k}=(c_{x}^{k},c_{y}^{k},w^{k},h^{k})$ described by the bounding box center $(c^{k}_{x},c^{k}_{y})$, width $w^{k}$ and height $h^{k}$, and its category information $\mathbf{a}^{k}$ describing the element type, such as text or image. In our considerations, the elements on the canvas do not possess an inherent order, which is why the canvas is considered a set of elements. Layout generation can be divided into various tasks depending on the conditioning assumptions. For example, in the unconditional generation setting, it is assumed that no prior knowledge is available, while for the refinement task, a complete but coarse layout is given. Since Flow-Matching requires the input data to be continuous, we convert the categorical information $\mathbf{a}^{k}$ into a continuous embedding, denoted as $\tilde{\mathbf{a}}^{k}$, using $B$ Analog Bits [7]. To learn the flow, we define a data sample $\mathbf{x}\in\mathbb{R}^{N_{max}\cdot(4+B)}$, which represents a layout as the concatenation of all its elements $\mathbf{g}^{k}$ and $\tilde{\mathbf{a}}^{k}$, and pad the remaining dimensions if the number of elements is smaller than $N_{max}$, the maximum number of elements in a single layout across the dataset.
@@ 4.2 Training
An overview of the training procedure for the type-conditioned scenario is illustrated in Fig. 2 and summarized in Algorithm 1 for the unconditional case. In the first step, we randomly sample a layout $\mathbf{x}_{0}$ from a prior distribution $p_{0}$, e.g. a Gaussian distribution. This sample will typically not be well-aligned and some elements might not be located on the canvas. In addition, we also randomly select a time $t$, which determines the intermediate sample $\mathbf{x}_{t}$ on the flow trajectory used during training. For LayoutFlow, we train on a simple linear trajectory following [44, 32]. Therefore, the intermediate sample is simply obtained by linearly interpolating between the ground truth layout $\mathbf{x}_{1}$ and the initial sample $\mathbf{x}_{0}$. The same trajectory is not used again during training, since $\mathbf{x}_{0}$ and $t$ change with every iteration.
The network $f_{\theta}$ is trained to output the conditional vector field, corresponding to the derivative of the trajectory as defined in Eq. 1. For the linear trajectory case, this derivative is simply a constant, specifically the difference between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$. Intuitively, the network learns to output a direction pointing toward a data sample, which can be used during sampling to move the initial data samples toward a better prediction. While we train our network using linear trajectories, this does not mean that the model will also produce straight trajectories during sampling. As mentioned in Eq. 4, the linear trajectory, which is derived by the conditional vector field in Eq. 4, acts as a proxy to learn the actual vector field in Eq. 3 defined by Flow Matching.
In our experiments, we found that using only the Mean Squared Error (MSE) as a loss function tends to lead to layouts that are not well aligned, which results in suboptimal perceptual quality. This observation is similar to the issue described in [5]. We hypothesize that this can be attributed to the fact that the MSE penalizes all mistakes within the same distance equally, regardless of direction, whereas the alignment between elements is usually only measured along the horizontal and vertical direction. As a result, the minimization objective does not fully align with the perceptual quality. To tackle this issue, we add an $L_{1}$ regularization on the geometrical part of the output $u^{\mathbf{g}}_{\theta}$ as the $L_{1}$ loss encourages sparsity of the error vector, which can also be interpreted as minimizing the error along the axis dimensions. Therefore, our final training loss can be written as
@@ 4.4 Network Architecture
Here, we describe the actual implementation of the network $f_{\theta}$ to efficiently handle the layout. We first embed the elements of the sampled intermediate layout $\mathbf{x}_{t}$ and then use a Transformer to predict the conditional vector field as depicted in Fig. 2. First, we separately embed the geometrical information $\mathbf{g}^{k}_{t}$ and the element type $\tilde{\mathbf{a}}^{k}_{t}$ using a linear layer. To inform the network about which parts of the input act as a condition, a conditioning mask is embedded and added to the geometrical and type embedding. The example in Fig. 2 illustrates this for the type-conditioned case, where the network is informed through the masks $\mathbf{m}_{type}$ and $\mathbf{m}_{pos}$ that $\tilde{\mathbf{a}}^{k}$ is taken from $\mathbf{x}_{0}$ while only the geometrical information $\mathbf{g}^{k}_{t}$ was taken from the training trajectory. In a subsequent step, we concatenate both embeddings and pass them through another linear layer to obtain a high-dimensional embedding $\mathbf{f}_{t}^{k}$ representing an element of the intermediate layout $\mathbf{x}_{t}$. Next, all element embeddings $\mathbf{f}_{t}^{k}$ are processed by a transformer encoder network together with the temporal information $t$ incorporated in each layer. Finally, the output of the Transformer is projected back to the same dimension as the input layout using a linear layer representing the vector field.
@@ 5.1 Experimental Settings
Datasets. Following previous methods, we evaluate the performance of our network on the RICO [9] and PubLayNet [49] datasets. RICO consists of over 66k User Interface (UI) layouts with 25 element categories, while PubLayNet comprises over 360k document layouts annotated with 5 different element types. We train our network using the dataset split described in [48, 23], which discards layouts containing more than 20 elements.
Implementation Details. Following LayoutDM and DLT, we employ a 4-layer Transformer network with 8 attention heads and a model dimension of 512 for a fair comparison. The model is trained using the AdamW optimizer with a learning rate of 0.0005. In addition to the MSE Flow Matching loss, we regularize the training by calculating the $L_{1}$ loss between the geometrical information as described in Eq. 6 with $\lambda=0.2$. For inference, we obtain the flow using a NeuralODE solver [6] using 100 steps with the Euler method.
@@
Comparison with Existing Approaches. To validate the performance of our proposed method, we compare it to state-of-the-art methods. Since there is no standard way of splitting the datasets, we need to retrain the models that do not follow the dataset split used in LayoutDiffusion [48] for a fair comparison. This fact limits us to only compare to methods that have made their code publicly available. Therefore, we compare our method against the two discrete diffusion models LayoutDM [4] and LayoutDiffusion [48] as well as the discrete-continuous diffusion model DLT [28]. Where applicable, we additionally compare LayoutFlow to the non-diffusion-based models LayoutTransformer [14], LayoutFormer++ [23], NDN-none [27] and RUITE [38].
Tasks. We test LayoutFlow and existing approaches on various common unconditional and conditional layout generation tasks. Un-Gen describes the layout generation task without any constraints. For the conditional scenario, we consider Gen-Type as assuming the element types as given, while Gen-TypeSize assumes element types and sizes are known. Lastly, we consider the Completion task which takes in a partial layout that is missing up to 20% of its elements, and the Refinement task following [48] with a standard deviation of 0.01.
@@ 5.2 Evaluation
Quantitative Evaluation. In Tab. 9, we report the results of our experiments comparing LayoutFlow with existing approaches for different layout generation tasks. In terms of FID, our model outperforms state-of-the-art methods across all three tasks, except for a close second place on the PubLayNet Un-Gen task, proving its strong capabilities to model the underlying sample distribution. In particular, the FID score of 2.37 produced by LayoutFlow on the unconditional generation for RICO almost matches the score obtained when comparing the validation with the test dataset. Note that for the conditional tasks, the FID score can become even lower than the validation dataset since the generated images share more similarities through conditioning. Regarding the geometrical metrics Alignment, Overlap, and mIoU, our method consistently outperforms LayoutDM and DLT, while usually producing values close to LayoutDiffusion, a significantly larger model with 85M parameters compared to around 15M used by LayoutFlow. A comparison between the quality and the speed of sampling is illustrated in Fig. 8, clearly showing that our method closes the gap between inference speed and performance. While non-diffusion models outperform diffusion models on some geometrical metrics, the drastically higher FID score implies that the generated results lack diversity.
@@
LayoutDiffusion 0.55 0.102 0.469 0.719 2.045 0.035 0.008 0.660
LayoutFlow (ours) 0.77 0.152 0.455 0.700 0.180 0.020 0.005 0.723
@@
Qualitative Evaluation. We provide some qualitative examples in Fig. 6 for the unconditional Fig. 7 as well as for the conditional generation task on the RICO dataset. More samples for the other tasks and PubLayNet can be found in the supplementary material. Overall, LayoutFlow shows a strong performance across all tasks producing visually pleasing results resembling real layouts.
@@
Figure 7: Condtional Generation Examples

### v2

@@ 1 Introduction
Motivated by this observation, we propose LayoutFlow, a generative framework based on Flow Matching [32, 33, 1, 44] that can produce high-quality layouts while requiring less time than previous methods. Flow Matching has recently been introduced as a powerful generative framework and demonstrated strong performance on various tasks, including image generation [34]. Intuitively, flow-based models try to learn a flow that moves samples from a base distribution to a target distribution defined by the training data. In this work, we investigate how to apply Flow Matching for layout generation and show that flows offer a more intuitive generation process from a geometrical interpretation compared to previous models.
@@
In the context of layout generation, the differences between diffusion-based and flow-based models become apparent when visualizing the trajectories created by the generation process, as illustrated in Fig. 1. The trajectories show how the sample from the base distribution is moved by the flow or backward diffusion process until it reaches the final layout prediction. While our LayoutFlow model produces smooth trajectories with short path lengths, diffusion-based models follow a noisy trajectory that constantly changes directions due to the added noise. As a result, diffusion models tend to require more sampling steps, i.e., longer inference time, than flow-based models. Altogether, we argue that learning how to move the initial sample straight toward the final prediction is much more natural than trying to do so under additional noise. Overall, our contributions can be summarized as follows:
@@
    On top of generating simpler trajectories, flow-based models generally also offer more choices than diffusion models, such as different prior distributions or training trajectories. We, therefore, extensively explore this additional flexibility offered by Flow Matching in the context of layout generation.
@@ 2.1 Layout Generation
Over the years, layout generation has been investigated using various methods. Early works started exploring generating layouts by minimizing an energy function based on pre-defined constraints [37, 36]. With the advent of generative machine learning techniques, however, research shifted to a more data-centric learning approach and started to use Variational Autoencoders (VAE) [47, 13, 2, 27], and Generative Adversarial Networks (GAN) [29, 30, 25]. While earlier works explored a variety of different architectures, such as Graph Convolutional Networks [27], Long-Short Term Memories (LSTM) [13, 24] or Transformers [47, 45, 14], most recent research has converged to using a Transformer-based architecture due to its flexibility and strong performance.
Diffusion-based Models. As diffusion models have proven to generate more diverse data and are generally easy to train, especially compared to GANs, most recent research has shifted towards exploring the diffusion process for layout generation tasks. Since layouts are represented by categorical data defining types of elements and continuous numerical values describing the element position and size, applying standard diffusion models is not straightforward. As a result, different approaches have incorporated a diffusion loss into their training process, which can be divided into continuous [18] and discrete diffusion models [3]. For the discrete case, the continuous coordinates are quantized and interpreted as different states along with the categorical element type [22, 20, 48, 15]. In this discrete scenario, the forward diffusion process is modeled as a random walk between discrete states that, in addition to the quantized states, also include a mask state, which removes elements from the canvas. While LayoutDM [22] and LDGM [20] rely on the mask-and-replace strategy proposed in [12], LayoutDiffusion [48] introduces a new mild forward process that is closer to the continuous process while still increasingly masking elements over time.
@@ 2.2 Flow Matching and Diffusion Models
Diffusion models, which were first proposed by Sohl-Dickenstein et al. [40] and later popularized by Ho et al. [18], can be formulated similarly to flow-based models. In fact, the term Flow Matching was inspired by the similar Score Matching loss [21] used by diffusion models. However, flows are characterized by an ODE, whereas the diffusion process can be formulated as an SDE with an additional stochastic component described by Brownian Motion [42]. As a result, diffusion models are restricted to a Gaussian as a base distribution, while flow-based models allow for more flexibility, which we also explore in our experiments. Notably, Song et al. [41] proposed DDIM, an alternative sampling method that makes it possible to sample using an ODE formulation on a model trained with an SDE forward process, effectively leading to smoother trajectories. In the context of layout generation, however, models have been using the SDE formulation following DDPM [18]. Our argument for introducing Flow Matching instead of just using DDIM for smoother trajectories lies in its increased flexibility and simplicity. Moreover, if the goal is to sample using an ODE, it seems more natural to also use an ODE during training instead of the more complex SDE used in diffusion.
@@ 3 Preliminary: Flow Matching
In this section, we first present the mathematical definition of a flow and then introduce how to train a neural network to estimate a flow via Flow Matching. The term flow in Flow Matching refers to a mapping between samples of two distributions. In general, flow-based models aim to estimate the flow between a known source probability distribution $p_{0}(x)$, e.g., a Gaussian distribution, and the typically more complex data distribution $p_{1}(x)$. Given a data point $x\in\mathbb{R}^{d}$, the flow $\phi:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ can be defined by the following Ordinary Differential Equation (ODE)
@@
However, this loss alone does not allow us to train the model as we usually do not have access to $p_{t}$ nor $v_{t}$. As an extension, Lipman et al. [32] have proposed a Conditional Flow Matching objective and is given as
@@
with $v_{t}(x|z)$ denoting a conditional vector field generating the conditional probability path $p_{t}(x|z)$ and $q(z)$ being a distribution over a latent conditioning variable $z$. By introducing the conditional formulation of the Flow Matching objective, training a neural network becomes possible by choosing and training with a conditional vector field and its associated probability path. For an overview of the various choices and more details on Conditional Flow Matching, we recommend [32, 44].
@@ 4.1 Data Representation
where we split each element into its geometrical information $\mathbf{g}^{k}=(c_{x}^{k},c_{y}^{k},w^{k},h^{k})$ described by the bounding box center $(c^{k}_{x},c^{k}_{y})$, width $w^{k}$ and height $h^{k}$, and its category information $\mathbf{a}^{k}$ describing the element type, such as text or image. In our considerations, the elements on the canvas do not possess an inherent order, which is why the canvas is considered a set of elements. Layout generation can be divided into various tasks depending on the conditioning assumptions. For example, in the unconditional generation setting, it is assumed that no prior knowledge is available, while for the refinement task, a complete but coarse layout is given. Since Flow-Matching requires the input data to be continuous, we convert the categorical information $\mathbf{a}^{k}$ into a continuous embedding, denoted as $\tilde{\mathbf{a}}^{k}$, using $B$ Analog Bits [7]. To learn the flow, we define a data sample $\mathbf{x}\in\mathbb{R}^{(4+B)\cdot N_{max}}$ as the concatenation of all the elements $\mathbf{g}^{k}$ and $\tilde{\mathbf{a}}^{k}$ in a single layout. If the number of elements is smaller than $N_{max}$, the maximum number of elements contained in a single layout across the entire dataset, we pad the remaining dimensions.
@@ 4.2 Training
An overview of the training procedure for the type-conditioned scenario is illustrated in Fig. 2 and summarized in Algorithm 1 for the unconditional case. In the first step, we randomly sample a layout $\mathbf{x}_{0}$ from a prior distribution $p_{0}$, e.g., a Gaussian distribution. This sample will typically not be well-aligned, and some elements might not be located on the canvas. In addition, we also randomly select a time $t$, which determines the intermediate sample $\mathbf{x}_{t}$ on the flow trajectory used during training. For LayoutFlow, we train on a simple linear trajectory following [44, 32]. Therefore, the intermediate sample is simply obtained by linearly interpolating between the ground truth layout $\mathbf{x}_{1}$ and the initial sample $\mathbf{x}_{0}$. The same trajectory is not used again during training since $\mathbf{x}_{0}$ and $t$ change with every iteration.
The network $f_{\theta}$ is trained to output the conditional vector field, corresponding to the derivative of the trajectory as defined in Eq. 1. This derivative is simply a constant for the linear trajectory case, specifically the difference between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$. Intuitively, the network learns to output a direction pointing toward a data sample, which can be used during sampling to move the initial data samples toward a better prediction. While we train our network using linear trajectories, this does not mean that the model will also produce straight trajectories during sampling. As mentioned in Eq. 4, the linear trajectory, which is derived by the conditional vector field in Eq. 4, acts as a proxy to learn the actual vector field in Eq. 3 defined by Flow Matching.
In our experiments, we found that using only the Mean Squared Error (MSE) as a loss function tends to lead to poorly aligned layouts, which results in suboptimal perceptual quality. This observation is similar to the issue described in [5]. We hypothesize that this can be attributed to the fact that the MSE penalizes all mistakes within the same distance equally, regardless of direction, whereas the alignment between elements is usually only measured along the horizontal and vertical direction. As a result, the minimization objective does not fully align with the perceptual quality. To tackle this issue, we add an $L_{1}$ regularization on the geometrical part of the output $u^{\mathbf{g}}_{\theta}$ as the $L_{1}$ loss encourages sparsity of the error vector, which can also be interpreted as minimizing the error along the axis dimensions. Therefore, our final training loss can be written as
@@ 4.4 Network Architecture
Here, we describe the actual implementation of the network $f_{\theta}$ to efficiently handle the layout. We first embed the elements of the sampled intermediate layout $\mathbf{x}_{t}$ and then use a Transformer to predict the conditional vector field as depicted in Fig. 2. First, we separately embed the geometrical information $\mathbf{g}^{k}_{t}$ and the element type $\tilde{\mathbf{a}}^{k}_{t}$ using a linear layer. A conditioning mask is embedded and added to the geometrical and type embedding to inform the network about which parts of the input act as a condition. The example in Fig. 2 illustrates this for the type-conditioned case, where the network is informed through the masks $\mathbf{m}_{type}$ and $\mathbf{m}_{pos}$ that $\tilde{\mathbf{a}}^{k}$ is taken from $\mathbf{x}_{1}$ while only the geometrical information $\mathbf{g}^{k}_{t}$ was taken from the training trajectory. In a subsequent step, we concatenate both embeddings and pass them through another linear layer to obtain a high-dimensional embedding $\mathbf{f}_{t}^{k}$ representing an element of the intermediate layout $\mathbf{x}_{t}$. Next, all element embeddings $\mathbf{f}_{t}^{k}$ are processed by a transformer encoder network together with the temporal information $t$ incorporated in each layer. Finally, the output of the Transformer is projected back to the same dimension as the input layout using a linear layer representing the vector field.
@@ 5.1 Experimental Settings
Datasets. We evaluate the performance of our network on the RICO [9] and PubLayNet [49] datasets, following previous methods. RICO consists of over 66k User Interface (UI) layouts with 25 element categories, while PubLayNet comprises over 360k document layouts annotated with 5 different element types. We train our network using the dataset split described in [48, 23], which discards layouts containing more than 20 elements.
Implementation Details. Similar to LayoutDM and DLT, we employ a 4-layer Transformer network with 8 attention heads and a model dimension of 512 for a fair comparison. The model is trained using the AdamW optimizer with a learning rate of 0.0005. In addition to the MSE Flow Matching loss, we regularize the training by calculating the $L_{1}$ loss between the geometrical information as described in Eq. 6 with $\lambda=0.2$. For inference, we obtain the flow using a NeuralODE solver [6] for 100 steps with the Euler method.
@@
Comparison with Existing Approaches. To validate the performance of our proposed method, we compare it to state-of-the-art techniques. Since there is no standard way of splitting the datasets, we need to retrain the models that do not follow the dataset split used in LayoutDiffusion [48] for a fair comparison. This fact limits us to only compare to methods that have made their code publicly available. Therefore, we compare our method against the two discrete diffusion models LayoutDM [4] and LayoutDiffusion [48] as well as the discrete-continuous diffusion model DLT [28]. Where applicable, we additionally compare LayoutFlow to the non-diffusion-based models LayoutTransformer [14], LayoutFormer++ [23], NDN-none [27] and RUITE [38].
Tasks. We test LayoutFlow and existing approaches on various common unconditional and conditional layout generation tasks. Un-Gen describes the layout generation task without any constraints. For the conditional scenario, we consider Gen-Type as assuming the element types as given, while Gen-TypeSize assumes element types and sizes are known. Lastly, we consider the Completion task, which takes in a partial layout that is missing up to 20% of its elements, and the Refinement task following [48] with a standard deviation of 0.01.
@@ 5.2 Evaluation
Quantitative Evaluation. In Tab. 10, we report the results of our experiments comparing LayoutFlow with existing approaches for different layout generation tasks. In terms of FID, our model outperforms state-of-the-art methods across all three tasks, except for a close second place on the PubLayNet Un-Gen task, proving its strong capabilities to model the underlying sample distribution. In particular, the FID score of 2.37 produced by LayoutFlow on the unconditional generation for RICO almost matches the score obtained when comparing the validation with the test dataset. Note that for the conditional tasks, the FID score can become even lower than the validation dataset since the generated images share more similarities through conditioning. Regarding the geometrical metrics Alignment, Overlap, and mIoU, our method consistently outperforms LayoutDM and DLT while usually producing values close to LayoutDiffusion, a significantly larger model with 85M parameters compared to around 15M used by LayoutFlow. A comparison between the quality and the sampling speed is illustrated in Fig. 8, clearly showing that our method closes the gap between inference speed and performance. While non-diffusion models outperform diffusion models on some geometrical metrics, the drastically higher FID score implies that the generated results lack diversity.
@@
LayoutDiffusion 0.55 0.102 0.469 0.719 2.05 0.035 0.008 0.660
LayoutFlow (ours) 0.77 0.152 0.455 0.700 0.18 0.020 0.005 0.723
@@
Qualitative Evaluation. We provide some qualitative examples in Fig. 6 for the unconditional Fig. 7 as well as for the conditional generation task on the RICO dataset. More samples for the other tasks and PubLayNet can be found in the supplementary material. Overall, LayoutFlow shows a strong performance across all tasks, producing visually pleasing results that resemble real layouts.
Limitations. We discuss the limitations of our proposed methods in the supplementary material.
@@
Figure 7: Conditional Generation Examples